# On Central Primitives for Quantum Cryptography with Classical Communication

### v1

@@ 9.1.1 $\mathsf{OWPuzz}$ imply next-bit pseudoentropy
The proof is exactly the same as the proofs of Lemma 3.6 and Theorem 3.15 from [VZ12], but with the use of the Min-Max theorem replaced by its quantum equivalent (Theorem 4.1 from [${}^{+}$17]). ∎
@@
Let $(\mathsf{Samp},\mathsf{Ver})$ be a $(\omega,\gamma)$-$\mathsf{OWPuzz}$ secure against all time $t$ adversaries with puzzle length $m(\lambda)$. Then for all $\epsilon>0$, $\mathsf{Samp}\to(k,s)$ has quantum $(t^{\prime},\epsilon/m)$ next-bit pseudoentropy at least
@@ 9.1.2 Next-bit pseudoentropy implies non-uniform EFID
The reductions for all three of these lemmas from [HRV10] are fully black-box, and so also hold in the quantum setting. ∎
@@
If, for some $c>0$, there exists a $(\mathsf{negl}(\lambda),1-\lambda^{-c})$ one-way puzzle $(\mathsf{Samp},\mathsf{Ver})$, then there exists an $\nu^{*}$-non-uniform EFID pair with $\nu^{*}\leq m+n$.
@@ 9.2 QEFID imply $\mathsf{OWPuzz}$
10 Random Input $\mathsf{OWPuzz}\leftrightarrow\mathsf{OWPuzz}$
@@ 11 Acknowledgments
We thank Yanyi Liu for insightful discussion. E. Goldin was supported by a National Science Foundation Graduate Research Fellowship.

### v2

@@ 9.1.1 $\mathsf{OWPuzz}$ imply next-bit pseudoentropy
The proof is exactly the same as the proofs of Lemma 3.6 and Theorem 3.15 from [VZ12], but with the use of the Min-Max theorem replaced by its quantum equivalent (Theorem 4.1 from [CCL+17]). ∎
@@
Let $(\mathsf{Samp},\mathsf{Ver})$ be a $(\omega,\gamma)$-$\mathsf{OWPuzz}$ with puzzle length $m(\lambda)$ secure against all time $t$ adversaries. Then for all $\epsilon>0$, $\mathsf{Samp}\to(k,s)$ has quantum $(t^{\prime},\epsilon/m)$ next-bit pseudoentropy at least
@@ 9.1.2 Next-bit pseudoentropy implies non-uniform EFID
The reductions for all three of these lemmas from [HRV10] are fully black-box (in a quantum friendly way), and also hold in the quantum setting. ∎
@@
If, for some $c>0$, there exists a $(\mathsf{negl}(\lambda),1-\lambda^{-c})$ one-way puzzle $(\mathsf{Samp},\mathsf{Ver})$, then there exists a $\nu^{*}$-non-uniform EFID pair with $\nu^{*}\leq m+n$.
@@ 9.2 QEFID imply $\mathsf{OWPuzz}$
10 Equivalence of $\mathsf{OWPuzz}$ to variants
10.1 Random Input $\mathsf{OWPuzz}\leftrightarrow\mathsf{OWPuzz}$
@@
10.2 Distributional $\mathsf{OWPuzz}\leftrightarrow\mathsf{OWPuzz}$
Here we show how to build $\mathsf{OWPuzz}$ from distributional $\mathsf{OWPuzz}$.
Definition 17.
A $\beta$ distributional one-way puzzle is a uniform QPT algorithm $\mathsf{Samp}(1^{\lambda})\to(k,s)$ which takes in a security parameter $\lambda$ and produces a key $k$ and a puzzle $s$ such that given a puzzle $s$, it is computationally infeasible to sample from the conditional distribution over keys. More formally, we require that for all non-uniform QPT algorithms $\mathcal{A}$, for all sufficiently large $\lambda$,
$\mathsf{Samp}(1^{\lambda})\to(k,s)$
$\Delta((k,s),(\mathcal{A},s))\geq\beta(\lambda)$
Theorem 33.
If, for some $c>0$, there exists a $\lambda^{-c}$ distributional one-way puzzle $\mathsf{Samp}$, then there exists a strong one-way puzzle.
This theorem follows directly from Pinsker’s inequality, which states
Theorem 34 (Pinsker’s inequality).
Let $P$ and $Q$ be any two probability distributions. Then
$\Delta(P,Q)\leq\sqrt{\frac{\ln 2}{2}KL(P||Q)}.$
Thus, if a puzzle is $\lambda^{-c}$ distributionally one-way, then we have that for all QPT $S$,
$KL(s,k||s,S(s))\geq\frac{12}{\ln 2}\Delta(s,k||s,S(k))^{2}\geq\lambda^{-2c}$
Reading the proof of Theorem 23, it is clear that the only property required of the one-way puzzle is that
$KL(s,k||s,S(s))\geq\frac{1}{\mathsf{poly}(\lambda)},$
and so applying the construction of Theorem 23 to $\mathsf{Samp}$ gives a strong one-way puzzle.
@@ 11 Acknowledgments
We thank Yanyi Liu for insightful discussion. We thank Bruno Cavalar for discussions concerning distributional one way puzzles. Kai-Min Chung was partially supported by the Air Force Office of Scientific Research under award number FA2386-23-1-4107 and NSTC QC project, under Grant no. NSTC 112-2119-M-001-006. E. Goldin was supported by a National Science Foundation Graduate Research Fellowship.